OpenStudy (anonymous):

What polynomial has roots of -5, -4, and 1?

OpenStudy (anonymous):

A. x3 - 8x2 - 11x + 20 B. x3 - x2 + 22x + 40 C. x3 + x2 - 22x - 40 D. x3 + 8x2 + 11x - 20

OpenStudy (anonymous):

I'll medal and fan help

OpenStudy (michele_laino):

for example, what is the root of this polynomial: (x+5)

OpenStudy (anonymous):

eek.. don't know. I suck that's why I'm asking for help.

OpenStudy (campbell_st):

do you know how to substitute into an equation?

OpenStudy (anonymous):

a little.

OpenStudy (anonymous):

a thrid order Polynomial can be written as: \[y = (x-a)(x-b)(x-c)\] where a, b and c are the roots of the polynomial. Have you ever seen it?

OpenStudy (campbell_st):

ok...the factor theorem uses substitution. It basically says... is a is a root then f(a) = 0

OpenStudy (anonymous):

I also put all the equations into mathway.. and they all didn't have the same roots.. maybe I typed one wrong.. idk

OpenStudy (anonymous):

?

OpenStudy (anonymous):

If you have an equation f(x) and suppose "a" is a root of this function, then f(a) = 0 That is, x=a will make function f(x) be zero

OpenStudy (anonymous):

you want to find a polymonial that have three root, -5,-4 and 1

OpenStudy (anonymous):

If we have three root, we must find a third order polynomial, since the order of the polynomial defines the number of root it have

OpenStudy (anonymous):

So, we need to find f(x) such that f(-5) = f(-4) = f(1) = 0

OpenStudy (anonymous):

From polynomial theory, we can write a polynomial like that: \[y = (x-a)(x-b)(x-c)\] where a,b and c are the roots

OpenStudy (anonymous):

Fine so far?

OpenStudy (anonymous):

Getting the hang out it

OpenStudy (anonymous):

now just plug the roots you have in the equation form: \[y = (x-(-5))(x-(-4))(x-1)\]\[y = (x+5)(x+4)(x-1)\]

OpenStudy (anonymous):

what would I do with it after

OpenStudy (anonymous):

make the distribution of factors